# Squeezed Quasinormal Modes from Nonlinear Gravitational Effects

### v1

@@ Abstract
We demonstrate that long-lived black hole quasi-normal modes become quantum mechanically squeezed due to nonlinear gravitational effects. We estimate the degree of squeezing for the fundamental mode of a Schwarzschild black hole quantitatively.
@@ I Introduction
Consideration of the interaction of radiation with quantum detectors requires some assumption, implicit or explicit, about the quantum nature of the radiation. Classical treatment of the radiation, amounting to the replacement of its field strength by c-numbers, corresponds implicitly to the assumption that it is embedded in the quantum framework as a coherent state (coherent state hypothesis). Squeezing takes us outside the domain of coherent states, and brings in characteristically quantum-mechanical effects. Since nonlinearities at the source or in propagation generically induce squeezing, and general relativity is an inherently nonlinear theory, some amount of squeezing is likely to be present in all forms of gravitational radiation. It is a challenging task, however, to make quantitative estimates for realistic, complex sources. Here we address this challenge for the radiation emitted during the “ringdown” following black hole formation.
Black hole quasi-normal modes Regge and Wheeler (1957); Zerilli (1970); Vishveshwara (1970); Berti et al. (2009); Kokkotas and Schmidt (1999); Berti et al. (2025) describe this stage of radiation. Although their amplitude is relatively weak, they are (relatively) tractable mathematically. Indeed, earlier works have estimated the nonlinear-gravitational effects systematically in the observable ringdown signal Tomita (1974); Tomita and Tajima (1976); Cheung et al. (2023); Gleiser et al. (1996, 2000); Mitman et al. (2023); Nakano and Ioka (2007); Ioka and Nakano (2007). Here we build upon these works to explore its implications for squeezing.
Expansion of Einstein’s field equations to the second order about a black hole spacetime,
@@
Above $G_{\mu\nu}^{(i)},~i=1,2$ refers to the Einstein tensor $G_{\mu\nu}$ expanded to appropriate order Tomita (1974); Tomita and Tajima (1976); Cheung et al. (2023); Gleiser et al. (1996, 2000); Mitman et al. (2023); Nakano and Ioka (2007); Ioka and Nakano (2007). In the Schwarzschild case, the first of the above equations leads to Regge-Wheeler Regge and Wheeler (1957), and Zerilli Zerilli (1970) equations for odd and even parity black hole quasi-normal modes, while the second equation leads to similar equations with a source term that is, up to derivatives, quadratic in the first order metric perturbations Cheung et al. (2023); Mitman et al. (2023); Nakano and Ioka (2007); Ioka and Nakano (2007). Recent works along this direction have noted that the coupled nature of wave equations at the second order imply higher harmonic generation of black hole quasinormal modes. There are good prospects to observe second harmonic generation sourced by lower order quasinormal modes, since the expected conversion ratio for lowest-order modes of non-rotating black holes is approximately, $\frac{|A_{4,4}(2\omega)|}{|A_{2,2}(\omega)^{2}|}\sim 0.15$ Cheung et al. (2023); Mitman et al. (2023); Nakano and Ioka (2007); Ioka and Nakano (2007). Perturbation theory for rotating (Kerr) spacetimes, notably developed by Teukolsky Teukolsky (1972) is more involved, but here too higher-harmonic quasinormal modes arise Ma and Yang (2024).
@@
In what follows, we present a general framework inspired by these recent developments to describe potential probes for quantum effects in weakly damped (long-lived) quasinormal modes. The suggestion to use the nonlinearities in black hole ringdown as a probe for quantum gravity effects, based on the analogy to second harmonic generation in quantum optics, was also made in Ref. Guerreiro (2023). Here we provide a systematic approach. After deriving an effective interaction Hamiltonian, we estimate the degree of squeezing that it leads to quantitatively. In interesting limits we can relate our problem to a parametrically excited oscillator, a context in which squeezing has been quantified both theoretically and experimentally Chaturvedi et al. (2002); Mallet et al. (2011); Murch et al. (2013); Wu et al. (1987); Adamyan et al. (2015); Manikandan and Qvarfort (2023). Our work complements recent discussions of how the coherent state hypothesis for gravitational radiation might be tested in practice with LIGO-type interferometers Parikh et al. (2020, 2021a, 2021b), and especially resonant mass detectors Tobar et al. (2024); Manikandan and Wilczek (2025a, ); Tobar et al. (2024); Manikandan and Wilczek (2025b); Shenderov et al. (2024).
Although we invoke most specifically black hole perturbation theory applied to the Schwarzschild background, the central ideas of our analysis are more widely applicable. Observed black hole events in LIGO correspond to the rotating (Kerr) black holes, where the lifetime of the fundamental quasinormal mode can be substantial, and in principle very large, becoming infinite in extremal limits Berti et al. (2009); Detweiler (1980); Yang et al. (2013a, b). While such modes are also difficult to excite, they provide limiting cases where our approach to estimating quantum effects, based on assuming long-lived modes, becomes exact. We defer a detailed analysis of such generalizations to future work.
II A simple theoretical model to describe quantum effects in quasinormal modes

### v2

@@ Abstract
We estimate the degree of squeezing possible in gravitational waves due to nonlinear gravitational effects in the weakly perturbative regime. Using the predicted amplitude ratios for higher harmonic generation in the ringdown phase of a black hole merger event, we estimate the relevant degree of squeezing produced by a Schwarzschild singularity to be of the order of one percent.
@@ I Introduction
Consideration of the interaction of radiation with quantum detectors requires some assumption, implicit or explicit, about the quantum nature of the radiation. Classical treatment of the radiation, amounting to the replacement of its field strength by c-numbers, corresponds implicitly to the assumption that it is embedded in the quantum framework as a coherent state (coherent state hypothesis). Squeezing takes us outside the domain of coherent states, and brings in characteristically quantum-mechanical effects. Since nonlinearities at the source or in propagation generically induce squeezing, and general relativity is an inherently nonlinear theory, some amount of squeezing is likely to be present in all forms of gravitational radiation. It is a challenging task, however, to make quantitative estimates for realistic, complex sources. Here we address this challenge by developing a simple theoretical framework that takes the amplitude ratio of higher order gravitational waves produced by nonlinear gravitational self-interactions in the background field as the input. While our model is generically applicable, we make quantitative estimates for the degree of squeezing using the amplitude ratios for radiation emitted during the “ringdown” following black hole formation, which gives a conservative estimate for the relevant degree of nonlinearity.
Black hole quasi-normal modes Regge and Wheeler (1957); Zerilli (1970); Vishveshwara (1970); Berti et al. (2009); Kokkotas and Schmidt (1999); Berti et al. (2025) describe the ringdown stage of radiation. Although their amplitude is relatively weak, they are (relatively) tractable mathematically. Indeed, earlier works have estimated the nonlinear-gravitational effects systematically in the observable ringdown signal Tomita (1974); Tomita and Tajima (1976); Cheung et al. (2023); Gleiser et al. (1996, 2000); Mitman et al. (2023); Nakano and Ioka (2007); Ioka and Nakano (2007). Expansion of Einstein’s field equations to the second order about a black hole spacetime,
@@
Above $G_{\mu\nu}^{(i)},~i=1,2$ refers to the Einstein tensor $G_{\mu\nu}$ expanded to appropriate order Tomita (1974); Tomita and Tajima (1976); Cheung et al. (2023); Gleiser et al. (1996, 2000); Mitman et al. (2023); Nakano and Ioka (2007); Ioka and Nakano (2007). In the Schwarzschild case, the first of the above equations leads to Regge-Wheeler Regge and Wheeler (1957), and Zerilli Zerilli (1970) equations for odd and even parity black hole quasi-normal modes, while the second equation leads to similar equations with a source term that is, up to derivatives, quadratic in the first order metric perturbations Cheung et al. (2023); Mitman et al. (2023); Nakano and Ioka (2007); Ioka and Nakano (2007).
@@
The suggestion to use the nonlinearities in black hole ringdown as a probe for quantum gravity effects, based on the analogy to second harmonic generation in quantum optics, was also made in Ref. Guerreiro (2023). Following a systematic analysis using simple dimensional arguments, we show that the degree of squeezing for such quadratic gravitational self-interactions turns out to be the square of the amplitude ratios. Using these, we predict an order of one percent squeezing of the phase-space quadratures in this tractable perturbative regime. In interesting limits we also relate our problem to a parametrically excited oscillator, a context in which squeezing has been quantified both theoretically and experimentally Chaturvedi et al. (2002); Mallet et al. (2011); Murch et al. (2013); Wu et al. (1987); Adamyan et al. (2015); Manikandan and Qvarfort (2023); Das et al. (2025). Here again, our analysis shows that the estimated degree of squeezing is of the order of the square of the amplitude ratio between the nonlinearly produced component and the fundamental mode. Our work complements recent discussions of how the coherent state hypothesis for gravitational radiation might be tested in practice with LIGO-type interferometers Parikh et al. (2020, 2021a, 2021b), and especially resonant mass detectors Tobar et al. (2024); Manikandan and Wilczek (2025a, ); Tobar et al. (2024); Manikandan and Wilczek (2025b); Shenderov et al. (2024).
II Quantum mechanical description of nonlinear gravitational wave interactions
Gravitational radiation, owing to inherent nonlinearities of gravity, is expected to have some degree of squeezing. Estimating the degree of squeezing involved is the challenging task. However gravitational wave scenarios where the relevant degree of nonlinearity can be estimated systematically gives hope that the degree of squeezing involved is both non-trivial and predictable. Higher harmonic generation in quasinormal modes provide one such framework which gives a conservative estimate for the relevant degree of nonlinearity involved in terms of calculable amplitude ratios. Several recent works have noted that the coupled nature of wave equations at the second order imply higher harmonic generation of black hole driven quasinormal modes, with leading contributions to the $l=4,m=4$ second-order quasinormal modes sourced by the $l=2,m=2$ mode Tomita (1974); Tomita and Tajima (1976); Cheung et al. (2023); Gleiser et al. (1996, 2000); Mitman et al. (2023); Nakano and Ioka (2007); Ioka and Nakano (2007). The “quasinormal” aspect reflects the fact that the mode frequencies involved are complex.
In the forthcoming discussions, we consider a toy model inspired by these developments, where we treat the frequencies involved approximately real. Strictly speaking, this approximation is only valid to describe early-time behavior of long-lived quasinormal modes. Therefore much of what follows can be understood as a rigorous toy model for normal mode gravitational waves with nonlinear self-interactions. Nevertheless, we keep the mode labels from the quasinormal mode context in all our forthcoming discussions, since we eventually would like to estimate the degree of squeezing in terms of the relevant nonlinearities in a gravitational scenario, which we extract from the amplitude ratios predicted for higher harmonic generation in quasinormal modes. There are good prospects to observe second harmonic generation sourced by lower order quasinormal modes, since the expected conversion ratio for lowest-order modes of non-rotating black holes is approximately, $\frac{|h_{44q}|}{|h_{220}^{2}|}\sim 0.15$ Cheung et al. (2023); Mitman et al. (2023); Nakano and Ioka (2007); Ioka and Nakano (2007). Perturbation theory for rotating (Kerr) spacetimes, notably developed by Teukolsky Teukolsky (1972) is more involved, but here too higher-harmonic quasinormal modes arise Ma and Yang (2024). Since quasinormal modes are decaying modes, we expect that the amplitude ratios predicted here provide a conservative estimate for the relevant degree of nonlinearity in comparable normal mode settings as well.
To the first order, the $l=2,m=2$ fundamental mode obeys the following wave equation,
$\ddot{h}_{220}+\omega_{220}^{2}h_{220}=0.$ (3)
Up to second order, a temporal wave equation describing the second harmonic generation of modes will have the generic form
$\ddot{h}_{44q}+\omega_{44j}^{2}h_{44q}\approx\sqrt{2}\Lambda_{j}h_{220}^{2},$ (4)
where $\Lambda_{j}$ with $\Lambda_{j}]=\left[T^{-2}\right]$ is an effective coupling to be determined. A framework to derive this coupled wave equation in the quasinormal mode context is presented in Appendix B where some of the challenges facing a rigorous derivation are also outlined. The frequency $\omega_{44j}$ appearing in the above equation exemplifies that, in the context of quasinormal modes, the modes that get pushed to second harmonic are the $l=4,m=4$ modes fundamental modes, sourced by the $l=2,m=2$ mode to the leading order. When the self interaction $\Lambda_{j}$ is neglected, the equation correspond to a first order quasinormal mode with $l=4,m=4$ that exist in the original spectrum. Looking at the form of the wave equation, and using Fourier analysis, one can conclude that the forced second harmonic mode will actually oscillate at twice the fundamental mode frequency. Alternatively, assuming a solution $h_{220}(t)\approx a\text{Re}(e^{-i\omega_{220}t})=a\cos{(\omega_{220,R}t)}e^{-\omega_{220,I}t}\approx a\cos{(\omega_{220}t)}$, for the fundamental mode [where we have also neglected damping, $\omega_{220}\approx\text{Re}(\omega_{220})$], we obtain that the particular solution to the above equation has the following form,
$h_{44j}^{(p)}(t)=\frac{a^{2}\Lambda_{j}}{\sqrt{2}\omega_{44j}^{2}}+\frac{a^{2}\Lambda_{j}\cos(2\omega_{220}t)}{\sqrt{2}(\omega_{44j}^{2}-4\omega_{220}^{2})}.$ (5)
Note that the temporal part of the particular solution oscillates at frequency $\omega_{44}^{(p)}=2\omega_{220}$, which is the second harmonic generated mode due to the effective forcing by the source term.
To compute quantum effects in a meaningful way it is important to estimate the effective interaction Hamiltonian in the quantum description, considering gravitational wave frequencies that are approximately real. As mentioned, long-lived quasinormal modes can very well be approximated in this manner. The quadratically sourced wave equation, Eq. (4) can be derived from the lagrangean,
$L_{q}=\left[\frac{1}{2}\left(\dot{h}_{44q}\right)^{2}-\frac{\omega_{44j}^{2}}{2}\left(h_{44q}\right)^{2}+\sqrt{2}\Lambda_{j}h_{220}^{2}h_{44q}\right].$ (6)
To this, we should add the lagrangean for the fundamental mode for completeness, given by,
$L_{0}=\frac{1}{2}\dot{h}_{220}^{2}-\frac{\omega_{220}^{2}}{2}h_{220}^{2},$ (7)
which suggests that some sort of back-reaction to the fundamental mode is inevitable in the presence of nonlinear interactions. It also helps to note that all the $44j$ modes could get pushed to the second harmonic to different degrees, determined by the coupling strengths $\Lambda_{j}$, but we expect that the near-resonant lower-$j$ modes would be the ones that will get maximally pushed. We can therefore replace $\Lambda_{j}$ with an effective coupling $\Lambda$ in the forthcoming discussions.
The wave equation for the fundamental mode also modifies to account for the non-linear interactions. Variation of the total Lagrangian $L=L_{q}+L_{0}$ with respect to $\dot{h}_{220}$ yields,
$\ddot{h}_{220}+\left[\omega_{220}^{2}-2\sqrt{2}\Lambda h_{44q}\right]h_{220}=0.$ (8)
We see that the mode $h_{44q}$ modulates the frequency of the fundamental mode in a time-dependent manner. Since the temporal part of $h_{44q}$ oscillates at frequency $\omega_{p}=2\omega_{220}$, we can approximate it as $h_{44q}(t)\approx h_{0}^{(q)}\cos(2\omega_{220}t)$. With this, we note that the equation governing the fundamental mode is modified as,
where $\epsilon=\sqrt{2}\Lambda h_{0}^{(q)}$. This is Mathieu’s equation in the standard form that describes a parametrically driven oscillator with $\omega_{p}=2\omega_{220}$. Considering the corresponding quantum oscillator problem, we note that such a drive naturally squeezes the fundamental mode as it can be understood as time-dependent modulations of the quantum harmonic potential, which has a squeezing effect Chaturvedi et al. (2002); Mallet et al. (2011); Murch et al. (2013); Wu et al. (1987); Adamyan et al. (2015); Manikandan and Qvarfort (2023); Das et al. (2025). The corresponding degree of squeezing, to the leading order, increases linearly with time Manikandan and Qvarfort (2023); Das et al. (2025),
$r_{\text{sq}}(t)\approx\epsilon t/(2\omega_{220})=\Lambda h_{0}^{(p)}t/(\sqrt{2}\omega_{220}).$ (10)
An initial coherent state gets squeezed as,
$\langle(\Delta X)^{2}\rangle=\frac{1}{2}e^{-2r}=\frac{1}{2}(1-2r)\approx\frac{1}{2}-r.$ (11)
We estimate the degree of squeezing it predicts in Sec. III.
The overtone mode attaches a primary degree of freedom of the quantized radiation field well outside the black hole to the nonlinear driving term that excites it. Similar situations have been treated in quantum optics and quantum acoustics, where the feedback from such modes has been shown to induce squeezing in the primary mode. Informed and inspired by that work, in the remainder of this section we provide a more systematic estimate in the framework of quantum theory, through construction of an effective Hamiltonian.
II.1 Effective Hamiltonian and dimensional analysis
An alternative semi-quantitative approach to estimate the degree of squeezing can be obtained with the help of dimensional analysis. We begin by noticing that both Eq. (3) and Eq. (4) we started with are for the strain amplitudes which are dimensionless. In order to map to quantum harmonic oscillators in the corresponding quantum theory and squeezing at appropriate time scales of interest, it is important to understand what is the effective quadrupole moment of these oscillators in a corresponding quantum description. This can be achieved by promoting the total lagrangean to one that has the dimensions of energy. To this end, we note that the lagrangean must be multiplied by a quantity that has dimensions of a mass quadrupole moment, $\mu]=[ML^{2}]$:
$\mu L_{q}=\mu\left[\frac{1}{2}\left(\dot{h}_{44q}\right)^{2}-\frac{\omega_{44j}^{2}}{2}\left(h_{44q}\right)^{2}+\sqrt{2}\Lambda h_{220}^{2}h_{44q}\right],$ (12)
such that,
$\left[\mu L_{q}\right]=[ML^{2}T^{-2}],$ (13)
has the dimensions of energy. Given that gravitational radiation is quadrupolar in nature, and sources that create are time-varying quadrupole moments, we can predict the scaling of $\mu$ with fundamental constants, and the frequency of the mode to be,
$\mu]=\left[\frac{c^{5}}{G\omega^{3}}\right].$ (14)
To estimate $\mu$ more definitively, we note that the corresponding wave equation for the second harmonic mode have the following form,
$\mu\ddot{h}_{44q}+\mu\omega_{44j}^{2}h_{44q}\approx\sqrt{2}\Lambda\mu h_{220}^{2}.$ (15)
Here we can fix $\mu\sim\frac{c^{5}}{16\pi G\Omega^{3}}$ such that the kinetic term of the Hamiltonian that yields the inertial term of the above equation maps to the energy of the wave in the mode volume, $V=(c/\Omega)^{3}$ given by $(E/V)(c/\Omega)^{3}\sim\frac{c^{2}}{32\pi G}\langle\dot{h}^{2}\rangle(c/\Omega)^{3}$, averaged over cycles.
Now the corresponding quantum harmonic oscillator Hamiltonian can be quantized in the usual manner, which correspond to promoting the field quadratures to operators in the following standard procedure,
$h_{\omega_{j}}(t)\rightarrow\hat{h}_{\omega_{j}}(t)=\sqrt{\frac{\hbar}{2\mu_{j}\omega_{j}}}(A_{\omega_{j}}e^{-i\omega_{j}t}+A_{\omega_{j}}^{\dagger}e^{i\omega_{j}t})=\sqrt{\frac{8\pi G\hbar\omega_{j}^{2}}{c^{5}}}(A_{\omega_{j}}e^{-i\omega_{j}t}+A_{\omega_{j}}^{\dagger}e^{i\omega_{j}t})$ (16)
From the lagrangean, we can also estimate the corresponding interaction Hamiltonian. Making use of the simplified notation $h_{220}\equiv h_{\omega}$ and $h_{44q}\equiv h_{2\omega}$, we see that the interaction hamiltonian will have the form,
$H_{I}$ $\approx$ $-\sqrt{2}\Lambda\mu\left[\hat{h}_{\omega}(t)\right]^{2}\hat{h}_{2\omega}(t)$ (17)
$=$ $-\hbar\sqrt{\hbar}\Lambda\mu\omega^{3}\left(\frac{16\pi G}{c^{5}}\right)^{3/2}(A_{\omega}e^{-i\omega t}+A_{\omega}^{\dagger}e^{i\omega t})^{2}(A_{2\omega}e^{-i2\omega t}+A_{2\omega}^{\dagger}e^{i2\omega t})$
Since the mapping concerns the kinetic (or inertial, term), it also makes sense to evaluate $\mu$ at frequency $\Omega=\omega_{44q}$ for the coupling, since the particular solution $h_{44q}(t)$ is a (damped) sinusoidal wave at frequency $\omega_{44q}=2\omega_{220}$. Substituting the approximate value for $\mu$ evaluated at the frequency of the second Harmonic mode, $\mu\approx\frac{c^{5}}{16\pi G(2\omega)^{3}}$, we obtain,
$\lambda=\Lambda\sqrt{\frac{\pi G\hbar}{c^{5}}}=\Lambda\tau_{p}\sqrt{\pi}.$ (18)
where $\tau_{p}=\sqrt{\frac{G\hbar}{c^{5}}}$ is the Planck time.
To relate to some of the standard discussions of second harmonic generation in quantum optics, we can make a further canonical transformation,
and
so that the interaction Hamiltonian can be represented as,
II.2 Estimating the coupling $\lambda$ from phenomenology
Above we have reduced the problem to a single effective parameter $\lambda$. Here we describe a phenomenological model that allows for estimating the coupling $\lambda$ from observable amplitudes in rather generic scenarios. To this end, we note that the amplitude $\frac{\lambda}{2}$ is determined by the fraction of energy dissipated via second harmonic generation, which numerical general relativity provides good estimates, that the amplitude of the second harmonic mode can be $\sim 10\%$ to that of the fundamental mode in some scenarios Cheung et al. (2023); Mitman et al. (2023). In fact, we can take the semi-classical limit to treat the fundamental mode in a coherent state $A_{\omega}|\alpha_{\omega}\rangle=\alpha_{\omega}|\alpha_{\omega}\rangle$, and approximate the interaction Hamiltonian as,
Note that the above relation holds for the quantum mechanical amplitude of the second harmonic mode, not to be confused with its temporal mode profile. The temporal mode profile for the second harmonic mode oscillates at frequency $2\omega$ as indicated by the subscript. As expected from Eq. (5), the amplitude is proportional to the coupling $\Lambda$ as well as the square of the amplitude of the fundamental mode. One can also identify a timescale $t\sim 1/\omega$ that can be chosen such that the amplitudes agree.
From Eq. (24), it seems appropriate that we could determine $\lambda$ at short times from simulations or actual data as,
III Degree of Squeezing
The interaction Hamiltonian in Eq. (22) describes second harmonic generation, which is a well-known method in quantum optics to generate and probe sub-Poissonian states of the radiation field Mandel (1982). We can extend the analysis to the gravitational context as well, and proceed to estimate the degree of squeezing. To this end, we can use the amplitude ratios predicted for the ringdown phase to estimate the relevant nonlinearity in a weakly perturbative regime of quantum gravity. Note that the Hamiltonian from Eq. (22) leads to the following set of equations that describe second harmonic generation,
that suggests that squeezing of the sub-Poissonian nature Mandel (1982). We can also use a perturbatively exact analysis based on Refs. Ou (1994); Li and Kumar (1994); Suhara et al. (1996), revisited in Appendix. C. For coherent state initial conditions for the fundamental mode and second harmonic mode in vacuum, where $\langle\hat{X}^{2}_{j\omega}(0)\rangle=\langle\hat{P}^{2}_{j\omega}(0)\rangle=\frac{1}{2}$, one obtains that $\langle\hat{X}^{2}_{j\omega}(\tau)\rangle=\langle\hat{X}^{2}_{j\omega}(0)\rangle S_{j\omega X},~~\text{and}~~\langle\hat{P}^{2}_{j\omega}(\tau)\rangle=\langle\hat{P}^{2}_{j\omega}(0)\rangle S_{j\omega P},~~j=1,2.$ Here, $\tau=\frac{1}{\sqrt{2}}\lambda|\langle A_{\omega}(0)\rangle|t=\frac{1}{\sqrt{2}}\lambda|\alpha_{\omega}(0)|t$ and,
$S_{\omega X}$ $=$ $(1-\tau\tanh\tau)^{2}\text{sech}^{2}\tau+2\tanh^{2}\tau\text{sech}^{2}\tau,$
$S_{2\omega X}$ $=$ $\frac{1}{2}(\tanh\tau+\tau\text{sech}^{2}\tau)^{2}+\text{sech}^{4}\tau,$
$S_{\omega P}$ $=$ $\frac{1}{2}(\sinh\tau+\tau\text{sech}\tau)^{2}+\text{sech}^{2}\tau,$
$S_{2\omega P}$ $=$ $2\tanh^{2}\tau+(1-\tau\tanh\tau)^{2},$ (29)
which reduces to Eq. (28) at small $\tau$. For $\tau\gg 1$, one obtains that, $S_{\omega X}\rightarrow 4\tau^{2}e^{-2\tau}$, $S_{2\omega X}\rightarrow 1/2$, $S_{\omega P}\rightarrow e^{2\tau}/8$, and $S_{2\omega P}\rightarrow\tau^{2}$ Ou (1994); Li and Kumar (1994); Suhara et al. (1996). This suggests substantial squeezing can in principle be generated through this mechanism, especially an exponential enhancement in the squeezing of the $\hat{P}$ quadrature of the fundamental mode.
$|\tau|=\frac{1}{\sqrt{2}}\lambda|\alpha_{\omega}(0)|t\approx\sqrt{\frac{16\pi G\hbar\omega^{2}}{c^{5}}}\left|\frac{h_{2\omega}}{h_{\omega}^{2}}\right||\alpha_{\omega}(0)|,$ (30)
$|\tau|=\sqrt{\frac{16\pi G\hbar\omega^{2}}{c^{5}}}\left|\frac{h_{2\omega}}{h_{\omega}^{2}}\right|\sqrt{\frac{c^{5}}{32\pi G\hbar\omega^{2}}}h_{\omega}\approx\frac{1}{\sqrt{2}}\left|\frac{h_{2\omega}}{h_{\omega}^{2}}\right|h_{\omega}\approx 0.1h_{\omega},$ (31)
since $\left|\frac{h_{2\omega}}{h_{\omega}^{2}}\right|\sim 0.15$. While the amplitude of the gravitational radiation is substantially small at our detectors, $h_{\omega}^{D}\approx 10^{-22}$, at the source it can be substantially high, and can be of the order of unity, suggesting that $|\tau|\approx 0.1$. Assuming that the fundamental mode is in a coherent state, we are also in the regime of validity of Eq. (28),
$\langle(\Delta\hat{X}_{\omega})^{2}\rangle$ $\approx$ $\frac{1}{2}-\lambda^{2}t^{2}|\alpha_{\omega}|^{2}/4=\frac{1}{2}\left(1-|\tau|^{2}\right)=0.495,~~\text{and}$
$\langle(\Delta\hat{P}_{\omega})^{2}\rangle$ $\approx$ $\frac{1}{2}+\lambda^{2}t^{2}|\alpha_{\omega}|^{2}/4=\frac{1}{2}\left(1+|\tau|^{2}\right)=0.505.$ (32)
Based on the mapping to the parametrically excited oscillator, we can also make some upper-bound for the degree of squeezing using Eq. (10), assuming the time $t$ is approximately the time it takes for the second harmonic amplitude to go from 0 to $h_{2\omega}$, and keeping the amplitude $h_{0}^{(p)}$ to be the maximum achievable amplitude in this time. We find that,
$\lambda t=\Lambda t\sqrt{\frac{\pi G\hbar}{c^{5}}}\implies\Lambda t=\sqrt{\frac{c^{5}}{\pi G\hbar}}\lambda t=2\frac{h_{2\omega}}{h_{\omega}^{2}}\sqrt{\frac{8\pi G\hbar\omega^{2}}{c^{5}}}\sqrt{\frac{c^{5}}{\pi G\hbar}}=4\sqrt{2}\omega\frac{h_{2\omega}}{h_{\omega}^{2}}$ (33)
This yields $r=\Lambda h_{0}^{(p)}t/(\sqrt{2}\omega)=4\frac{h_{2\omega}^{2}}{h_{\omega}^{2}}$. The actual achievable squeezing will be smaller than this, given that the drive amplitude $h_{0}^{(p)}$ gradually increases from 0 to $h_{2\omega}$. The more careful estimate from Eq. (32) yields,
$r=|\tau|^{2}/2=\frac{1}{4}\frac{h_{2\omega}^{2}}{h_{\omega}^{2}}.$ (34)
We note that up to a numerical pre-factor the degree of squeezing estimated agrees with our estimate based on mapping to a parametric oscillator. Probably the parametric oscillator analysis overestimates the degree of squeezing: it assumes that that the second harmonic mode of predicted amplitude is available for the whole duration, while in practice it is only being generated due to non-linear interactions.
For reasons of concreteness and simplicity, we focused on a scenario where the fundamental mode begins in a coherent state. The merger dynamics may well prepare them in an acoherent quantum state, however. We defer a detailed analysis of this to future work.
IV Discussion
We have estimated the degree of squeezing generated in gravitational waves from non-linear gravitational self-interactions. The estimates are made using predicted amplitude ratios from the perturbative regime following the merger event. For this scenario, where only weak nonlinearities are involved, we find that the predicted higher harmonic generation of quasinormal modes suggests a $\sim 1\%$ squeezing. As mentioned immediately above, our simplifying assumptions about initial conditions are not unlikely to be overly pessimistic. More generally, earlier stages of radiation close to merger, where nonlinearities are stronger, can be expected to yield larger effects. Recent work suggests possibilities for testing these ideas experimentally Manikandan and Wilczek (2025a, , b); Parikh et al. (2020, 2021b, 2021a); Tobar et al. (2024); Shenderov et al. (2024).
Note added—. A paper recently posted to the arXiv reaches similar conclusions to ours in a different, but closely related problem involving gravitational radiation Kanno et al. (2025).
V Acknowledgments
FW is supported by the U.S. Department of Energy under grant Contract Number DE-SC0012567 and by the Swedish Research Council under Contract No. 335-2014-7424. SKM was supported in part by the Swedish Research Council under Contract No. 335-2014-7424, and the Wallenberg Initiative on Networks and Quantum Information (WINQ). SKM also acknowledges the support from the Department of Atomic Energy, Government of India, under Project Identification No. RTI4007. We thank Maulik Parikh and Igor Pikovski for stimulating discussions around these subjects.
Or alternatively, the number of gravitons,
Based on this, we can estimate the amplitude ratio,
Appendix B An effective approach to obtain the coupled wave equation in the temporal domain for black-hole quasinormal modes